OpenStudy (anonymous):

solve for x: 3x - 7 + 4x = 42

OpenStudy (anonymous):

combine like terms on the left side; add 7 to both sides; then divide by the coefficient of x to both sides.

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

I'm trying to get this. can you show work and all steps

OpenStudy (anonymous):

what you got when you combine 3x + 4x?

OpenStudy (anonymous):

3apples + 4apples = ???

OpenStudy (anonymous):

7x

OpenStudy (anonymous):

ok.. what you did is combine like terms on the left side. now your equation looks like this: 7x - 7 = 42

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the next step is to add 7 to both sides of the equal sign... what u got now?

OpenStudy (anonymous):

14 = 49

OpenStudy (anonymous):

the right side is correct.. but the left side isn't...

OpenStudy (anonymous):

here's the left side

OpenStudy (anonymous):

7x-7+7

OpenStudy (anonymous):

so the left gots to be 0.

OpenStudy (anonymous):

7apples - 7 + 7 = ???

OpenStudy (anonymous):

7

OpenStudy (anonymous):

lemme put it this way..

OpenStudy (anonymous):

using associative property...

OpenStudy (anonymous):

7x - 7 + 7 7x + (-7 + 7) can you simplify what's in the parenthesis?

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

so it should be 0.

OpenStudy (anonymous):

If you owe 7 pennies and somebody gives you 7 pennies. How many pennies do you have?

OpenStudy (anonymous):

none

OpenStudy (anonymous):

Yes 0!

OpenStudy (anonymous):

thats what I said befor but I was erong

OpenStudy (anonymous):

But you still have 7x on the left side of your = sign

OpenStudy (anonymous):

what did you have on the right side of the = sign?

OpenStudy (anonymous):

7x=?

OpenStudy (anonymous):

49

OpenStudy (anonymous):

7x=49 7x is like saying 7 times x divide both sides by 7 now

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

can 7/ into 49

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its 7 x=7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wow I did it.. thank you very much

OpenStudy (anonymous):

No problem! Good job!

OpenStudy (anonymous):

how about this one? 4x - 5 = x + 13

OpenStudy (anonymous):

You need to get x on only one side of the = sign

OpenStudy (anonymous):

start by subtracting x from both sides. What will that leave you with?

OpenStudy (anonymous):

so I should have 3x

OpenStudy (anonymous):

Yep... 3x.....what?

OpenStudy (anonymous):

3x -5 = 13

OpenStudy (anonymous):

Good....since 5 is negative you need to add 5 to both sides now

OpenStudy (anonymous):

3x = 18

OpenStudy (anonymous):

yep...what do you think you should do now to figure out x?

OpenStudy (anonymous):

/ 3 on each side

OpenStudy (anonymous):

Yep! You're good to go now!

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

your awesome, I'm glad you took the time to teach me a few things

OpenStudy (anonymous):

Thanks! Hope it helped!

OpenStudy (anonymous):

it sure did, but I have another problem I did not understand

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

7( x + 2) + 3 ( x - 3) = 25

OpenStudy (anonymous):

Okay...do you know how to distribute?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

So lets just work with the first part of the equation: 7(x+2)

OpenStudy (anonymous):

You want to "distribute" the 7 into the parenthesis

OpenStudy (anonymous):

7x +2

OpenStudy (anonymous):

7x is right, but what about the 2?

OpenStudy (anonymous):

or should it be 7x + 14

OpenStudy (anonymous):

Yes! 7x + 14!

OpenStudy (anonymous):

and I do the same with 3( x - 3)

OpenStudy (anonymous):

Now lets work with the second part of the equation: 3(x-3)

OpenStudy (anonymous):

Yes...do the same

OpenStudy (anonymous):

7x + 14 + 3x - 9 = 25

OpenStudy (anonymous):

Yes! 7x +14 + 3x - 9= 25 Let's clean up the left side now. if you combine the x factors and the numbers what do you get?

OpenStudy (anonymous):

10x + 14 - 9 = 25

OpenStudy (anonymous):

yes...but what is 14 - 9?

OpenStudy (anonymous):

lol, idk I though it had to be in there

OpenStudy (anonymous):

10x + 5 = 25

OpenStudy (anonymous):

Yep! How do we get 10x by itself on the left of the = sign?

OpenStudy (anonymous):

-5 on both sides

OpenStudy (anonymous):

Good! Do that :)

OpenStudy (anonymous):

ok so my answer should be x = 2

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

once again we did it

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

now it time for graphs for me

OpenStudy (anonymous):

Good luck! I really dislike graphing. :)

OpenStudy (anonymous):

Gotta go

OpenStudy (anonymous):

ok thanks later